OpenStudy (anonymous):

Which of the following is a factor of 32x3 + 4?

OpenStudy (ganpat):

32x3 + 4 = 0 32x3 = -4 x3 = -4/32 = -1/8 x = cube root (-1/8) x = cube root (i2 * 1/8) x = sq rt i * 1/2.. Hope u get this.. :D

OpenStudy (anonymous):

Thats not an answer choice o:

OpenStudy (ganpat):

what r options ??

OpenStudy (anonymous):

I got the answer lol. I posted the same question with the answers but thanks (: